A tropical fish tank at an aquarium is 4 feet high and has a volume of 500ft^3. The length of the tank is twice the width. What is the width of the tank to the nearest tenth of a foot?
Answer by lwsshak3(11628)

A tropical fish tank at an aquarium is 4 feet high and has a volume of 500ft^3. The length of the tank is twice the width. What is the width of the tank to the nearest tenth of a foot?
***
let x=width of tank
2x=length of tank
volume=length*width*height
2x*x*4=500
8x^2=500
x^2=500/8=62.5
x=√62.5
x=7.9
width of the tank=7.9 ft